# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 6.2.  Switching beween $n$ and $q_{T}-n$
Let $x_{\textnormal{left}}^{*}=:x_{\textnormal{left}}^{*}(m,n)$ be defined as before, and define $x_{\textnormal{left}}^{*}(m,q_{T}-n)$ to be the analogous number for the pair $(m,q_{T}-n)$, i.e., $x_{\textnormal{left}}^{*}(m,q_{T}-n)$ minimises $\langle x\alpha\rangle_{\textnormal{left}}$ in the range $q_{T}-n-m+1,q_{T}-n+m-1]$. Define $x_{\textnormal{right}}^{*}(m,n)$ and $x_{\textnormal{right}}^{*}(m,q_{T}-n)$ analogously. Then we have
$x_{\textnormal{left}}^{*}(m,n)>n\iff x_{\textnormal{right}}^{*}(m,q_{T}-n)<q_{T}-n.$
@@
$\langle-x_{\textnormal{left}}^{*}\alpha\rangle_{\textnormal{right}}=\langle x_{\textnormal{left}}^{*}\alpha\rangle_{\textnormal{left}}=\min_{n-m+1\leq x\leq n+m-1}\langle x\alpha\rangle_{\textnormal{left}}=\min_{-(n+m-1)\leq x\leq-(n-m+1)}\langle x\alpha\rangle_{\textnormal{right}}.$
@@
$\langle(q_{T}-x_{\textnormal{left}}^{*})\alpha\rangle_{\textnormal{right}}=\max_{q_{T}-n-m+1\leq x\leq q_{T}-n+m-1}\langle x\alpha\rangle_{\textnormal{right}}.$
This means that $q_{T}-x_{\textnormal{left}}^{*}=x_{\textnormal{right}}^{*}(m,q_{T}-n)$, and so $x_{\textnormal{left}}^{*}>n\iff x_{\textnormal{right}}^{*}(m,q_{T}-n)<q_{T}-n$. ∎
6.3.  Technical lemmas for the case $x^{*}<n$
If $x^{*}<n$, then we do not have $x(\ell)=x^{*}$ for all $\ell\in[0,m-1]$. However, it is not hard to see that $x(\ell)=x^{*}$ for a certain range of $\ell$’s (which, admittedly, might be only the single integer $\ell=n-m+1$). We can say more about the precise values of $x(\ell)$ in adjacent ranges. We start with the special case where $m$ is particularly small. Recall that we always imply that either $(x^{*},x,\langle\cdot\rangle)=(x_{\textnormal{left}}^{*},x_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{left}})$ or $(x^{*},x,\langle\cdot\rangle)=(x_{\textnormal{right}}^{*},x_{\textnormal{right}},\langle\cdot\rangle_{\textnormal{right}})$.
@@ 6.3.  Technical lemmas for the case $x^{*}<n$
Let $2\leq m\leq n$ and assume $m<q_{1}$ (or $m<q_{2}$ if $\alpha>1/2$). Assume $x^{*}<n$. Then $x(m-1+x^{*}-n)=x^{*}$ and $x(m+x^{*}-n)=x^{*}+1$. In particular, $f$ is not bijective.
@@
Since $X(m-1+x^{*}-n)=[x^{*},x^{*}+m-1]$, we have $x(m-1+x^{*}-n)=x^{*}$ by definition of $x$ and $x^{*}$. Now consider $X(m+x^{*}-n)=[x^{*}+1,x^{*}+m]$. Let $x^{\prime}:=\min\{x>x^{*}\colon\langle x\alpha\rangle\}$. Since $x^{*}$ minimises $\langle x\rangle$ in the range [nm+1,n+m1][n-m+1,n+m-1], we have $x^{\prime}>n+m-1\geq x^{*}+m$. Since we are assuming $m<q_{1}$ (or $m<q_{2}$ if $\alpha>1/2$), Lemma 5.14 implies that $x(m+x^{*}-n)=x^{*}+1$. Thus, $x(m-x^{*}-n)-x(m-1+x^{*}-n)=1=(m-x^{*}-n)-(m-1+x^{*}-n)$, and so $f$ is not bijective by Lemma 6.3. ∎
Now we determine the value of $x(\ell)$ in certain ranges of $\ell$ in the case that $m$ is not very small. This is probably the most technical part of the paper; after that we will be able to prove (non-)balancedness in various cases.
@@
Assume $2\leq m\leq n$ and $q_{M-1}<m\leq q_{M}$ with $m\geq q_{2}$ (or $m\geq q_{1}$ if $\alpha<1/2$). Moreover, assume that $x^{*}<n$. Then we have
$x(\ell)=x^{*}\quad\iff\quad\ell\in[0,m-1+x^{*}-n]=:R_{0}.$
If $\langle q_{M-1}\alpha\rangle<1/2$, then
$x(\ell)=x^{*}+q_{M-1}$
$\iff\quad\ell\in[m+x^{*}-n,\min\{m+x^{*}-n-1+q_{M-1},m-1\}]=:R_{1}.$
If $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M}$, then
$x(\ell)=x^{*}+q_{M}\quad\iff\quad\ell\in[m+x^{*}-n,m-1]=:R_{1}^{\prime}.$
If $\langle q_{M}\alpha\rangle<1/2$ and $q_{M-2}+aq_{M-1}\leq m<q_{M-2}+(a+1)q_{M-1}$ with $0\leq a\leq a_{M}-1$, and where we assume that $M-2\geq 0$ (or $M-2\geq 1$ if $\alpha>1/2$), then
$x(\ell)=x^{*}+q_{M-2}+aq_{M-1}$
$\qquad\text{for all }\ell\in[m+x^{*}-n,\min\{x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1,m-1\}]=:R_{1}^{\prime\prime};$
$x(\ell)=x^{*}+q_{M-2}+(a+1)q_{M-1}$
$\qquad\text{for all }\ell\in[x^{*}-n+q_{M-2}+(a+1)q_{M-1},$
$\qquad\qquad\qquad\qquad\min\{x^{*}-n+q_{M-2}+(a+2)q_{M-1}-1,m-1\}]=:R_{2}^{\prime\prime}.$
Note that the range $R_{2}^{\prime\prime}$ is empty if $\min\{x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1,m-1\}=m-1$.
@@
By the definitions of $x^{*}$ and the function $x$, we clearly have $x(\ell)=x^{*}$ if and only if $x^{*}\in X(\ell)=[n+\ell-m+1,n+\ell]$. Under our assumption $x^{*}<n$, this happens exactly for $\ell\in[0,m-1+x^{*}-n]$. This settles the statement regarding the range $R_{0}$. Now we go through all the cases from the statement of the lemma.
Case 1: $\langle q_{M-1}\alpha\rangle<1/2$.
First, note that $x^{*}+q_{M-1}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[x^{*}-n+q_{M-1},m+x^{*}-n-1+q_{M-1}]$. Since $x^{*}-n+q_{M-1}\leq m+x^{*}-n$, we have that indeed $x^{*}+q_{M-1}\in X(\ell)$ for all $\ell\in R_{1}=[m+x^{*}-n,\min\{m+x^{*}-n-1+q_{M-1},m-1\}]$ (and not for any larger $\ell$). Now we only need to show that $x^{*}+q_{M-1}$ indeed minimises $\langle x\alpha\rangle$ in each range $X(\ell)$ for $\ell\in R_{1}$. In other words, we need to show that
(6.3) $\langle(x^{*}+q_{M-1})\rangle=\min\{\langle x\alpha\rangle\colon x\in[x^{*}+1,\min\{m+x^{*}-1+q_{M-1},n+m-1\}]\}.$
From the definition of $x^{*}$ it follows that
$x^{\prime}:=\min\{x>x^{*}\colon\langle x\alpha\rangle<\langle x^{*}\alpha\rangle\}>n+m-1.$
Moreover, note that $m+x^{*}-1+q_{M-1}<x^{*}+q_{M-1}+q_{M}$. Therefore, we get (6.3) directly from Lemma 5.15 with $a=0$. This settles the statement regarding the range $R_{1}$.
Case 2: $\langle q_{M}\alpha\rangle<1/2$ and $m=q_{M}$
First, note that $x^{*}+q_{M}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[x^{*}-n+q_{M},m+x^{*}-n-1+q_{M}]=[m+x^{*}-n,2m+x^{*}-n-1]$. Since $2m+x^{*}-n-1\geq m$, we have in fact $x^{*}+q_{M}\in X(\ell)$ for all $\ell\in R_{1}^{\prime}=[m+x^{*}-n,m-1]$. Now we only need to show that $x^{*}+q_{M}$ indeed minimises $\langle x\alpha\rangle$ in each range $X(\ell)$ for $\ell\in R_{1}^{\prime}$. Since the largest element in all these ranges is $n+m-1\leq x^{*}+2m-2=x^{*}+2q_{M}-2<x^{*}+q_{M}+q_{M+1}$, this follows again directly from Lemma 5.15 with $a=0$.
Case 3: $\langle q_{M}\alpha\rangle<1/2$ and $q_{M-2}+aq_{M-1}\leq m<q_{M-2}+(a+1)q_{M-1}$ with $0\leq a\leq a_{M}-1$.
First, note that $x^{*}+q_{M-2}+aq_{M-1}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[x^{*}-n+q_{M-2}+aq_{M-1},m+x^{*}-n-1+q_{M-2}+aq_{M-1}]$. On the one hand, we have $x^{*}-n+q_{M-2}+aq_{M-1}\leq m+x^{*}-n$. On the other hand, recall that in the lemma we assume $m>q_{M-1}$, and so $m+x^{*}-n-1+q_{M-2}+aq_{M-1}\geq x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1$. Thus, we have indeed $x^{*}+q_{M-2}+aq_{M-1}\in X(\ell)$ for $\ell\in R_{1}^{\prime\prime}=[m+x^{*}-n,\min\{x^{*}-n+q_{M-2}+(a+1)q_{M-1}-1,m-1\}]$. We need to check that $x^{*}+q_{M-2}+aq_{M-1}$ minimises $\langle x\alpha\rangle$ in the range $x^{*}+1,x^{*}+q_{M-2}+(a+1)q_{M-1}-1]$. Indeed, this follows directly from Lemma 5.15.
In the case $x^{*}<n$ we can now use the above lemma to characterise $f$ being bijective in terms of the shape of $x^{*}$ and $m$.

### v2

@@ 6.2.  Switching between $n$ and $q_{T}-n$
Let $\widehat{x}_{\textnormal{left}}(m,n)=\widehat{x}_{\textnormal{left}}$ be defined as before, and define $\widehat{x}_{\textnormal{right}}(m,q_{T}-n)$ in an analogous way, i.e., $\widehat{x}_{\textnormal{right}}(m,q_{T}-n)$ minimises $\langle x\alpha\rangle_{\textnormal{right}}$ in the range $q_{T}-n-m+1,q_{T}-n+m-1]$. Then for sufficiently large $T$ we have
$\widehat{x}_{\textnormal{left}}(m,n)>n\iff\widehat{x}_{\textnormal{right}}(m,q_{T}-n)<q_{T}-n.$
@@
$\langle-\widehat{x}_{\textnormal{left}}\alpha\rangle_{\textnormal{right}}=\langle\widehat{x}_{\textnormal{left}}\alpha\rangle_{\textnormal{left}}=\min_{n-m+1\leq x\leq n+m-1}\langle x\alpha\rangle_{\textnormal{left}}=\min_{-(n+m-1)\leq x\leq-(n-m+1)}\langle x\alpha\rangle_{\textnormal{right}}.$
@@
$\langle(q_{T}-\widehat{x}_{\textnormal{left}})\alpha\rangle_{\textnormal{right}}=\max_{q_{T}-n-m+1\leq x\leq q_{T}-n+m-1}\langle x\alpha\rangle_{\textnormal{right}}.$
This means that $q_{T}-\widehat{x}_{\textnormal{left}}=\widehat{x}_{\textnormal{right}}(m,q_{T}-n)$, and so $\widehat{x}_{\textnormal{left}}>n\iff\widehat{x}_{\textnormal{right}}(m,q_{T}-n)<q_{T}-n$. ∎
6.3.  Technical lemmas for the case $\widehat{x}_{*}<n$
If $\widehat{x}_{*}<n$, then we do not have $x_{*}(\ell)=\widehat{x}_{*}$ for all $\ell\in[0,m-1]$. However, it is not hard to see that $x_{*}(\ell)=\widehat{x}_{*}$ for a certain range of $\ell$’s (which, admittedly, might be only the single integer $\ell=n-m+1$). We can say more about the precise values of $x_{*}(\ell)$ in adjacent ranges. This is probably the most technical part of the paper; after that we will be able to prove (non-)balancedness in various cases.
Recall that we always implicitly assume either $(f_{*},x_{*},\langle\cdot\rangle_{*},\widehat{x}_{*})=(f_{\textnormal{left}},x_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{left}},\widehat{x}_{\textnormal{left}})$ or $(f_{*},x_{*},\langle\cdot\rangle_{*},\widehat{x}_{*})=(f_{\textnormal{right}},x_{\textnormal{right}},\langle\cdot\rangle_{\textnormal{right}},\widehat{x}_{\textnormal{right}})$. For technical reasons, we assume $m\geq q_{1}$; the cases where $m$ is very small are actually quite easy and we will deal with them separately in Section 6.5.
@@ 6.3.  Technical lemmas for the case $\widehat{x}_{*}<n$
Assume $2\leq q_{1}\leq m\leq n$ and $q_{M-1}<m\leq q_{M}$. Moreover, assume that $\widehat{x}_{*}<n$. Then we have
$x_{*}(\ell)=\widehat{x}_{*}\quad\iff\quad\ell\in[0,m-1+\widehat{x}_{*}-n]=:R_{0}.$
If $\langle q_{M-1}\alpha\rangle_{*}<1/2$, then
$x_{*}(\ell)=\widehat{x}_{*}+q_{M-1}$
$\iff\quad\ell\in[m+\widehat{x}_{*}-n,\min\{m+\widehat{x}_{*}-n-1+q_{M-1},m-1\}]=:R_{1}.$
If $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M}$, then
$x_{*}(\ell)=\widehat{x}_{*}+q_{M}\quad\iff\quad\ell\in[m+\widehat{x}_{*}-n,m-1]=:R_{1}^{\prime}.$
If $\langle q_{M}\alpha\rangle_{*}<1/2$ and $q_{M-2}+aq_{M-1}\leq m<q_{M-2}+(a+1)q_{M-1}$ with $0\leq a\leq a_{M}-1$, then
$x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+aq_{M-1}$
$\qquad\text{for all }\ell\in[m+\widehat{x}_{*}-n,\min\{\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1,m-1\}]=:R_{1}^{\prime\prime};$
$x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+(a+1)q_{M-1}$
$\qquad\text{for all }\ell\in[\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1},$
$\qquad\qquad\qquad\qquad\min\{\widehat{x}_{*}-n+q_{M-2}+(a+2)q_{M-1}-1,m-1\}]=:R_{2}^{\prime\prime}.$
Note that the range $R_{2}^{\prime\prime}$ is empty if and only if $m-1\leq\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1$.
@@
By the definitions of $\widehat{x}_{*}$ and $x_{*}(\cdot)$, we clearly have $x_{*}(\ell)=\widehat{x}_{*}$ if and only if $\widehat{x}_{*}\in X(\ell)=[n+\ell-m+1,n+\ell]$. Under our assumption $\widehat{x}_{*}<n$, this happens exactly for $\ell\in[0,m-1+\widehat{x}_{*}-n]$. This settles the statement regarding the range $R_{0}$. Now we go through all the cases from the statement of the lemma.
Case 1: $\langle q_{M-1}\alpha\rangle_{*}<1/2$.
First, note that $\widehat{x}_{*}+q_{M-1}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[\widehat{x}_{*}-n+q_{M-1},m+\widehat{x}_{*}-n-1+q_{M-1}]$. Since $\widehat{x}_{*}-n+q_{M-1}\leq m+\widehat{x}_{*}-n$, we have indeed $\widehat{x}_{*}+q_{M-1}\in X(\ell)$ for all $\ell\in R_{1}=[m+\widehat{x}_{*}-n,\min\{m+\widehat{x}_{*}-n-1+q_{M-1},m-1\}]$ (and not for any larger $\ell$). Now we only need to show that $\widehat{x}_{*}+q_{M-1}$ indeed minimises $\langle x\alpha\rangle_{*}$ in each range $X(\ell)$ for $\ell\in R_{1}$. In other words, we need to show that
(6.3) $\langle(\widehat{x}_{*}+q_{M-1})\alpha\rangle_{*}=\min\{\langle x\alpha\rangle_{*}\colon x\in[\widehat{x}_{*}+1,\min\{m+\widehat{x}_{*}-1+q_{M-1},n+m-1\}]\}.$
From the definition of $\widehat{x}_{*}$ it follows that
$x^{\prime}:=\min\{x>\widehat{x}_{*}\colon\langle x\alpha\rangle_{*}<\langle\widehat{x}_{*}\alpha\rangle_{*}\}>n+m-1.$
Moreover, note that $m+\widehat{x}_{*}-1+q_{M-1}<\widehat{x}_{*}+q_{M-1}+q_{M}$. Therefore, we get (6.3) directly from Lemma 5.14 with $a=0$. This settles the statement regarding the range $R_{1}$.
Case 2: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M}$
First, note that $\widehat{x}_{*}+q_{M}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[\widehat{x}_{*}-n+q_{M},m+\widehat{x}_{*}-n-1+q_{M}]=[m+\widehat{x}_{*}-n,2m+\widehat{x}_{*}-n-1]$. Since $2m+\widehat{x}_{*}-n-1\geq m$, we have in fact $\widehat{x}_{*}+q_{M}\in X(\ell)$ for all $\ell\in R_{1}^{\prime}=[m+\widehat{x}_{*}-n,m-1]$. Now we only need to show that $\widehat{x}_{*}+q_{M}$ indeed minimises $\langle x\alpha\rangle_{*}$ in each range $X(\ell)$ for $\ell\in R_{1}^{\prime}$. Since the largest element in all these ranges is $n+m-1\leq\widehat{x}_{*}+2m-2=\widehat{x}_{*}+2q_{M}-2<\widehat{x}_{*}+q_{M}+q_{M+1}$, this follows again directly from Lemma 5.14 with $a=0$.
Case 3: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $q_{M-2}+aq_{M-1}\leq m<q_{M-2}+(a+1)q_{M-1}$ with $0\leq a\leq a_{M}-1$.
First, note that $\widehat{x}_{*}+q_{M-2}+aq_{M-1}\in X(\ell)=[n+\ell-m+1,n+\ell]$ if and only if $\ell\in[\widehat{x}_{*}-n+q_{M-2}+aq_{M-1},m+\widehat{x}_{*}-n-1+q_{M-2}+aq_{M-1}]$. On the one hand, we have $\widehat{x}_{*}-n+q_{M-2}+aq_{M-1}\leq m+\widehat{x}_{*}-n$. On the other hand, recall that in the lemma we assume $m>q_{M-1}$, and so $m+\widehat{x}_{*}-n-1+q_{M-2}+aq_{M-1}\geq\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1$. Thus, we have indeed $\widehat{x}_{*}+q_{M-2}+aq_{M-1}\in X(\ell)$ for $\ell\in R_{1}^{\prime\prime}=[m+\widehat{x}_{*}-n,\min\{\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1,m-1\}]$. We need to check that $\widehat{x}_{*}+q_{M-2}+aq_{M-1}$ minimises $\langle x\alpha\rangle_{*}$ in the range $\widehat{x}_{*}+1,\widehat{x}_{*}+q_{M-2}+(a+1)q_{M-1}-1]$. Indeed, this follows directly from Lemma 5.14.
In the case $\widehat{x}_{*}<n$ we can now use the above lemma to characterise $f_{*}$ being bijective in terms of the shape of $\widehat{x}_{*}$ and $m$. Again, we assume $m\geq q_{1}$.
@@
Assume $2\leq q_{1}\leq m\leq n$ and $q_{M-1}<m\leq q_{M}$. Moreover, assume that $\widehat{x}_{*}<n$. Then $f_{*}$ is bijective in exactly the three following cases:
  • $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M}$;
  • $\langle q_{M}\alpha\rangle_{*}<1/2$, $m=q_{M-2}+aq_{M-1}$ for some $1\leq a\leq a_{M}-1$ and $M-2\geq 0$, and $\widehat{x}_{*}\geq n-q_{M-1}$;
  • $\langle q_{M}\alpha\rangle_{*}<1/2$, $q_{M-2}+aq_{M-1}<m<q_{M-2}+(a+1)q_{M-1}$ for some $0\leq a\leq a_{M}-1$ and $\widehat{x}_{*}=n-q_{M-1}$.
@@
We distinguish between five cases, according to whether $\langle q_{M}\alpha\rangle_{*}<1/2$ or $\langle q_{M-1}\alpha\rangle_{*}<1/2$, and some extra conditions.
Case 1: $\langle q_{M-1}\alpha\rangle_{*}<1/2$.
Then, by Lemma 6.8, we have $x_{*}(\ell)=\widehat{x}_{*}$ for $\ell\in R_{0}$ and $x_{*}(\ell)=\widehat{x}_{*}+q_{M-1}$ for $\ell\in R_{1}$. Since $m-1\geq q_{M-1}$ and $m+\widehat{x}_{*}-n-1+q_{M-1}\geq q_{M-1}$, the range $R_{0}\cup R_{1}$ contains at least $q_{M-1}+1$ consecutive integers. Therefore, there exist $\ell\in R_{0}$ and $\ell^{\prime}\in R_{1}$ such that $\ell^{\prime}-\ell=q_{M-1}$. For these $\ell,\ell^{\prime}$ we have $x_{*}(\ell^{\prime})-x_{*}(\ell)=q_{M-1}=\ell^{\prime}$-$\ell$, and so $f_{*}$ is not bijective by Lemma 6.3.
Case 2: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M}$.
By Lemma 6.8, we have either $x_{*}(\ell)=\widehat{x}_{*}$ or $x_{*}(\ell)=\widehat{x}_{*}+q_{M}$ for all $\ell\in R_{0}\cup R_{1}^{\prime}=[0,m-1]=[0,q_{M}-1]$. Therefore, it is clear that we cannot have $\ell-\ell^{\prime}=x_{*}(\ell)-x_{*}(\ell^{\prime})$ for $\ell\neq\ell^{\prime}\in[0,m-1]$, and so $f_{*}$ is bijective by Lemma 6.3.
Case 3: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$ and $\widehat{x}_{*}\geq n-q_{M-1}$.
Then, by Lemma 6.8, we have $x_{*}(\ell)=\widehat{x}_{*}$ for $\ell\in R_{0}$ and $x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+aq_{M-1}$ for $\ell\in R_{1}^{\prime\prime}$. Moreover, since we are assuming $\widehat{x}_{*}\geq n-q_{M-1}$, we have
$\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1\geq q_{M-2}+aq_{M-1}-1=m-1.$
In other words, $R_{0}\cup R_{1}^{\prime\prime}=[0,m-1]$, and so for all $\ell\in[0,m-1]=[0,q_{M-2}+aq_{M-1}-1]$ either $x_{*}(\ell)=\widehat{x}_{*}$ or $x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+aq_{M-1}$. As in Case 2, it is clear that $f_{*}$ is bijective.
Case 4: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $m=q_{M-2}+aq_{M-1}$ with $1\leq a\leq a_{M}-1$, but now $\widehat{x}_{*}<n-q_{M-1}$.
Then
$\widehat{x}_{*}-n+q_{M-2}+(a+1)q_{M-1}-1<m-1,$
and so there is at least one $\ell\in R_{2}^{\prime\prime}$. Now since the range $R_{1}^{\prime\prime}$ contains exactly $q_{M-1}$ integers (note that $m=q_{M-2}+aq_{M-1}$), the range $R_{1}^{\prime\prime}\cup R_{2}^{\prime\prime}$ contains at least $q_{M-1}+1$ consecutive integers. Moreover, Lemma 6.8 says that $x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+aq_{M-1}$ for $\ell\in R_{1}^{\prime\prime}$ and $x_{*}(\ell)=\widehat{x}_{*}+q_{M-2}+(a+1)q_{M-1}$ for $\ell\in R_{2}^{\prime\prime}$. Thus, there must exist $\ell\in R_{1}^{\prime\prime}$ and $\ell^{\prime}\in R_{2}^{\prime\prime}$ with $\ell^{\prime}-\ell=q_{M-1}=x_{*}(\ell^{\prime})-x_{*}(\ell)$, and so $f_{*}$ is not bijective.
Case 5: $\langle q_{M}\alpha\rangle_{*}<1/2$ and $q_{M-2}+aq_{M-1}<m<q_{M-2}+(a+1)q_{M-1}$ for some $0\leq a\leq a_{M}-1$.
If $\widehat{x}_{*}>n-q_{M-1}$, then one can check that $R_{0}\cup R_{1}^{\prime\prime}\supseteq[0,q_{M-2}+aq_{M-1}]$. Therefore, there must exist $\ell\in R_{0}$ and $\ell^{\prime}\in R_{1}^{\prime\prime}$ with $\ell^{\prime}-\ell=q_{M-2}+aq_{M-1}=x_{*}(\ell^{\prime})-x_{*}(\ell)$. Thus, $f_{*}$ is not bijective.
If $\widehat{x}_{*}<n-q_{M-1}$, then one can check that the range $R_{1}^{\prime\prime}\cup R_{2}^{\prime\prime}$ contains at least $q_{M-1}+1$ consecutive integers, and that $R_{1}^{\prime\prime}$, $R_{2}^{\prime\prime}$ are non-empty, and so, as in Case 4, $f_{*}$ is not bijective.
If $\widehat{x}_{*}=n-q_{M-1}$, then the first term in the minimum in the upper bound of $R_{2}^{\prime\prime}$ is
$\widehat{x}_{*}-n+q_{M-2}+(a+2)q_{M-1}-1=q_{M+2}+(a+1)q_{M-1}-1\geq m-1,$
so $R_{0}\cup R_{1}^{\prime\prime}\cup R_{2}^{\prime\prime}=[0,m-1]$. Also, one can check that $R_{0}\cup R_{1}^{\prime\prime}\not\supseteq[0,q_{M-2}+aq_{M-1}]$, and that $R_{1}^{\prime\prime}\cup R_{2}^{\prime\prime}$ contains at most $q_{M-1}$ consecutive integers. Thus, $x_{*}(\ell^{\prime})-x_{*}(\ell)\neq\ell^{\prime}$-$\ell$ for all $\ell\in R_{0}$ and $\ell^{\prime}\in R_{1}^{\prime\prime}$, as well as for all $\ell\in R_{1}^{\prime\prime}$ and $\ell^{\prime}\in R_{2}^{\prime\prime}$. It is also easy to see that $x_{*}(\ell^{\prime})-x_{*}(\ell)\neq\ell^{\prime}$-$\ell$ for all $\ell\in R_{0}$ and $\ell\in R_{2}^{\prime\prime}$ because $m-1<q_{M-2}+(a+1)q_{M-1}$. Thus, $f_{*}$ is bijective. ∎
6.4.  The cases where $\widehat{x}_{*}\neq n$
First, we show that the third special case from Lemma 6.9 actually corresponds to the case where $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$.